OpenStudy (zaxoanl):

the label on a hershey's chocolate bar claims 210 Calories per 37 gram serving. How much would you expect the temperature of the water to rise if 1.32 grams of chocolate were burned and the mass of the water was 150 grams

OpenStudy (zaxoanl):

please help i dont know how to set up the equations

OpenStudy (aaronq):

The Calories given in the specifications of food is not the thermal energy upon combustion but through cellular respiration lol... ignoring that.. we can use: \(q=m*C_p*\Delta T\) We use the water to monitor the combustion reaction (because the energy released by it is absorbed by the water). \(q_{absorbed}=m_{water}*C_p*\Delta T\) solve for the change in temperature: \(\Delta T=\dfrac{q_{absorbed}}{C_p*m_{water}}\)

OpenStudy (anonymous):

Note that food "calories" are actually kilocalories. (210 kcal) / (37 g) = 5.6756 kcal/g = 23747 J/g (23747 J/g) x (1.32 g) / (4.184 J/g·°C) / (150 g) = 50 °C change